Circular No. 7676

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 O2 (NEAT) Regarding the announcement of this comet on IAUC 7673, the obser-

vations on July 25 were made at Haleakala, while those on July 29 were made at Palomar. Additional astrometry and the following very uncertain parabolic orbital elements (from 17 observations, July 25–Aug. 1) and an ephemeris appear on MPEC 2001-P01:

T = 2000 May 12.560 TT  ω = 314.^{o}165
Ω = 326.634 }2000.0
q = 6.81841 AU  i = 101.120

SUPERNOVA 2001dk IN UGC 913 A. Gal-Yam and O. Shemmer, Tel Aviv University, report: “A spec-

trum (range 400–780 nm) of SN 2001dk (cf. IAUC 7675) was obtained with the Wise Observatory 1-m telescope (+ FOSC spectrograph) on July 31 by J. Dann. Underlying light from the host galaxy prevents positive identification of this object. The spectrum displays strong and relatively narrow (FWHM < 2000 km/s) hydrogen Balmer lines of Hα and Hβ, and its overall shape is similar to the spectrum of type-IIn SN 1994ak several weeks past maximum light (Filippenko 1997, Ann. Rev. A.Ap. 35, 309). A doublet around 672 nm may be identified with [S II] 671.6- and 673.1-nm.”

SUPERNOVA 2001dl IN UGC 11725 M. Armstrong, Rolvenden, Kent, England, reports that an unfiltered

CCD image taken by himself on July 30.994 UT with a 0.30-m reflector shows SN 2001dl (cf. IAUC 7675) at mag 16.4. An earlier image taken on July 28.003 shows a possible hint of the supernova at mag ∼ 18.5–19.0.

COMET C/2001 A2 (LINEAR) Visual m_{1} estimates: July 13.44 UT, 4.4 (C. S. Morris, Fillmore, CA,

10×50 binoculars); 14.95, 4.9 (M. Lehký, Hradec Králové, Czech Republic, naked eye); 18.04, 5.8 (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars); 20.94, 6.2 (R. J. Bouma, Groningen, The Netherlands, 7×50 binoculars); 22.95, 6.3 (W. Hasubick, Buchloe, Germany, 10×50 binoculars); 27.29, 6.7 (C. E. Spratt, Victoria, BC, 7×50 binoculars); 29.96, 7.2 (A. Trebacz, Niepolomice, Poland, 10×50 binoculars).

2001 August 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green